### v2

# Majorana fermions in Kitaev chains side-coupled to normal metals

Abstract
Majorana fermions, exotic particles with potential applications in quantum computing, have garnered significant interest in condensed matter physics. The Kitaev model serves as a fundamental framework for investigating the emergence of Majorana fermions in one-dimensional systems. We explore the intriguing question of whether Majorana fermions can arise in a NM side-coupled to a Kitaev chain (KC) in the topologically trivial phase. Our findings reveal affirmative evidence, further demonstrating that the KC, when in the topological phase, can induce additional Majorana fermions in the neighboring NM region. Through extensive parameter analysis, we uncover the potential for zero, one, or two pairs of Majorana fermions in a KC side-coupled to an NM. Additionally, we investigate the impact of magnetic flux on the system and calculate the winding number -a topological invariant - to characterize the phase transitions.
I Introduction
The Kitaev model stands as a cornerstone in the study of Majorana fermions (MFs) within condensed matter physics Kitaev (2001). This involves a p-wave superconducting term in the Hamiltonian of one-dimensional spinless normal metal (NM). When the superconducting pairing term is nonzero, the open system described by the Hamiltonian is in trivial phase without any MFs or in topological phase with a pair of MFs depending on whether $|\mu|<2t$ or $|\mu|>2t$ respectively, where $\mu$ is the chemical potential and $t$ is the hopping strength. Theoretical proposals for realizing MFs revolutionized the search for these exotic particles Lutchyn et al. (2010); Oreg et al. (2010). Experimental evidences for MFs were reported Mourik et al. (2012); Das et al. (2012). The model proposed by Kitaev has been experimentally realized recently Dvir et al. (2023). Extensions of MFs to periodically driven systems have been studied Thakurathi et al. (2013); Soori (2023).
The topologically trivial phase in Kitaev model is characterized by an insulating phase when the superconducting pairing term is switched off. In other words, when $\mu>2t$ ($\mu<-2t$) the band is completely filled (empty). Hence, turning on superconductivity does not result in MFs. The limit of $\mu<-2t$ corresponds to Bose-Einstein condensate (BEC) superconductors hosting strongly bound pairs when s-wave superconductivity is switched on in spin-half tight binding chain Chen et al. (2005); Randeria and Taylor (2014). Andreev reflection spectroscopy in such BEC superconductors is expected to show finite subgap conductance when the barrier separating the NM and superconductor is turned into a well Lewandowski et al. (2023). Physically this can be understood as due to the formation of Andreev bound state due to potential well at the interface. This means that though the BEC superconductors have no electrons in their NM state, the superconductivity develops in the neighboring metal region if the metal hosts electrons. A similar line of argument can be applied to Kitaev model in the trivial phase to phrase the following question. ‘Can MFs develop in a NM side-coupled to a Kitaev chain (KC) in the topologically trivial phase?’ (See Fig.1 for a schematic of the system). This paper answers this question in positive. Further, we find that when KC is in the topological phase, it possible that it can induce another pair of MFs in the neighboring NM region. Depending on the choice of parameters, there can be zero, one or two pairs of MFs in KC side-coupled to a NM. We consider the possibility of a magnetic flux piercing the region between KC and NM. We calculate winding number - a topological invariant for KC side-coupled to NM to characterize the phase of the system.
Figure 1: Schematic diagram of a Kitaev chain side-coupled to a normal metal. $t^{\prime}$ is the hopping amplitude between the two chains. $\phi$ quantifies the magnetic flux piercing through each placket.
II The Hamiltonian and the bulk spectrum
The Hamiltonian for the system under consideration is:
$H$ $=$ $\sum_{n=-\infty}^{\infty}[-t(c^{\dagger}_{n+1}c_{n}+{\rm h.c.})-\mu(c^{\dagger}_{n}c_{n}-\frac{1}{2})$ (1)
$+\Delta(c^{\dagger}_{n+1}c^{\dagger}_{n}+{\rm h.c.})-t(f^{\dagger}_{n+1}f_{n}e^{i\phi}+{\rm h.c.})$
$-t^{\prime}(c^{\dagger}_{n}f_{n}+{\rm h.c.})],$
where $c_{n}$ ($f_{n}$) annihilates an electron at site $n$ on KC (NM), $t$ is the hopping strength within KC, $te^{i\phi}$ is the complex hopping amplitude in the NM, $\mu$ is the chemical potential on KC, $\Delta$ is the pairing strength within KC, $t^{\prime}$ is the coupling between KC and NM. The phase factor $\phi$ corresponds to a magnetic flux threading through the rectangular placket between KC and NM nearest neighbor sites. We first convert the Hamiltonian in eq. (1) to momentum space:
$H_{k}$ $=$ $\begin{bmatrix}-2t\cos{k}-\mu&-t^{\prime}&2i\Delta\sin{k}&0\\ -t^{\prime}&-2t\cos{k^{\prime}}&0&0\\ 2i\Delta\sin{k}&0&2t\cos{k}+\mu&t^{\prime}\\ 0&0&t^{\prime}&2t\cos{k^{\prime}}\end{bmatrix}~{},$ (6)
where $k^{\prime}=k+\phi$.
Figure 2: Logarithm of the bulk gap in units of $t$ versus $t^{\prime}$ and $\mu$ for $\Delta=0.5t$ and $\phi=0$. The data points with $t^{\prime}=0$ are excluded.
Unless mentioned explicitly, $\phi=0$. Taking $\Delta=0.5t$, we numerically diagonalize the Hamiltonian in eq. (LABEL:eq:ham-k) and plot the logarithm of bulk gap in Fig. 2. We find that the bulk gap closes on two curves that intersect at $(\mu,t^{\prime})=(0,2)t$. In the limit $t^{\prime}=0$, NM and KC are decoupled and the gap closes trivially since the NM is not gapped. On three sides of the gap-closing lines, we plot the spectrum for a finite chain with $L=1500$ sites and zoom it around the zero energy in Fig. 3 for $\Delta=0.5t$. We find that for $(\mu,t^{\prime})=(\pm 5t,t)$, there is one pair of zero energy states, for $(\mu,t^{\prime})=(0,t)$ there are two pairs of zero energy states and for $(\mu,t^{\prime})=(0,4t)$, there are no zero energy states. Next, we need to check whether the zero energy states found so are MFs and characterize the different sides of the gap closing lines by a topological invariant.
Figure 3: Spectrum of finite coupled chains zoomed around zero energy for $L=1500$ sites on each chain on three sides of the gap closing line with $\Delta=0.5t$ for: (a) $\mu=\pm 5t$, $t^{\prime}=t$, (b) $\mu=0$, $t^{\prime}=t$, (c) $\mu=0$, $t^{\prime}=4t$. Number of pairs of zero energy states: (a) 1, (b) 2, (c) 0.
Figure 4: Winding number versus $t^{\prime}$ and $\mu$. Red dotted lines are the lines where $|\mu_{\sigma}|$ is $2t$. Red dotted lines match with the gap closing lines in Fig. 2. The data points with $t^{\prime}=0$ are excluded.
III Winding number
The finding that zero energy states which could possibly be MFs can develop in a hybrid made of two topologically trivial systems is interesting. To investigate whether the zero energy states have Majorana character, we start with the non-superconducting hybrid. The electron eigenenergies of the Hamiltonian eq. (LABEL:eq:ham-k) for $\Delta=0$, $\phi=0$ are $E=-\mu/2-2t\cos{k}+\sigma\sqrt{(\mu/2)^{2}+t^{\prime 2}}$, with $\sigma=\pm 1$. The respective eigenstates are $|\phi^{e}_{\sigma}\rangle=[\phi_{\sigma},0,0]^{T}$ for the electron sector, where $\phi_{\sigma}$ is the left eigenvector of $2\times 2$ block in the electron sector of the Hamiltonian eq. (LABEL:eq:ham-k). Similarly, the eigenenergies in the hole sector are $E=\mu/2+2t\cos{k}-\sigma\sqrt{(\mu/2)^{2}+t^{\prime 2}}$, with $|\phi^{h}_{\sigma}\rangle=[0,0,\phi_{\sigma}]^{T}$ being the eigenvectors. The two sets of bands corresponding to $\sigma=\pm 1$ can be thought of as bonding and anti-bonding orbitals of a two-level system. The effective chemical potential in the sector $\sigma$ is $\mu_{\sigma}=\mu/2-\sigma\sqrt{(\mu/2)^{2}+t^{\prime 2}}$. When $\mu_{\sigma}$ falls outside (inside) the range $(-2t,2t)$, the phase can be expected to be topologically trivial (nontrivial) in the sector $\sigma$.
Figure 5: Phase diagrams of the system for $\Delta=0.5t$, and different values of $\phi$. Winding number is plotted versus $t^{\prime}/t$ and $\mu/t$. (a) $\phi=0.25\pi$, (b) $\phi=0.49\pi$, (c) $\phi=0.5\pi$, (d) $\phi=0.51\pi$, (e) $\phi=0.75\pi$, (f) $\phi=\pi$. The data points with $t^{\prime}=0$ are excluded.
We project the $4\times 4$ Hamiltonian in eq. (LABEL:eq:ham-k) into the two sectors described by $\sigma=\pm 1$. In sector $\sigma$, the projected Hamiltonian is given by
$h^{\sigma}_{k}$ $=$ $\begin{bmatrix}\langle\phi^{e}_{\sigma}|H_{k}|\phi^{e}_{\sigma}\rangle&\langle\phi^{e}_{\sigma}|H_{k}|\phi^{h}_{\sigma}\rangle\\ \langle\phi^{h}_{\sigma}|H_{k}|\phi^{e}_{\sigma}\rangle&\langle\phi^{h}_{\sigma}|H_{k}|\phi^{h}_{\sigma}\rangle\end{bmatrix}~{}.$ (10)
Such a matrix has the form $h^{\sigma}_{k}=\epsilon_{k}\tau_{z}+\Delta_{k}\tau_{y}$, where $\epsilon_{k}$ and $\Delta_{k}$ are functions of $k$, and $\tau_{z}$ and $\tau_{y}$ are $2\times 2$ Pauli spin matrices. The winding number in sector $\sigma$ denoted by $w^{\sigma}$ is then given by the number of times $(\epsilon_{k},\Delta_{k})$ loops around $(0,0)$ when $k$ is taken from 00 to $2\pi$. The total winding number $w=w^{+}+w^{-}$ is a non negative integer which indicates the number of pairs of MFs in the open system. For each set of values of the parameters $(t^{\prime},\mu)$, we numerically find the winding $w$ and plot it in Fig. 4 for $\Delta=0.5t$. The red dotted lines indicate the values of the parameters for which $|\mu_{\sigma}|$ crosses $2t$. At the line separating the regions having $w=1$ and $w=2$, one of the two $\mu_{\sigma}$’s crosses ±$\pm 2t$. The same is true for the line separating the regions with winding numbers 00 and $1$.
In Fig. 4, we have excluded the data points for $t^{\prime}=0$, where the gap closes and the winding number is not well defined in one of the sectors. For small values of $t^{\prime}/t$, the gap is small in the sector $\sigma=1$. This sector corresponds to the electrons being populated mostly in the NM. Due to the small coupling to KC, superconductivity leaks into NM and a small gap opens up in this sector. This means that the decay length of MFs in the $\sigma=1$ sector gets larger as $t^{\prime}/t$ decreases. As $t^{\prime}/t$ becomes smaller, the length $L$ of the system needed to see two MFs becomes larger. This holds for the phase diagrams in Fig. 5.
IV Phase diagram for $\phi\neq 0$
The procedure of calculating the winding number described in the previous section holds good even when $\phi\neq 0$. We calculate the winding number for different values of $\phi$ in Fig. 5. We find that the phase diagram changes as $\phi$ is varied. As $\phi$ increases from 00 to $\pi/2$, the region corresponding to $w=2$ in the phase diagram shrinks. Interestingly, for the choice of $\phi\geq\pi/2$, there are only two phases: the one with single pair of MFs and another with no MFs. We can understand the limiting case of $\phi=\pi$ by the following argument. The dispersion relations for the two electron bands of the Hamiltonian are $E=-\mu/2\pm\sqrt{t^{\prime 2}+(\mu/2+2t\cos k)^{2}}$. The superconductivity induces MFs if the electron bands cross $E=0$. Of the two bands here, only one band can cross $E=0$. It can be shown with some algebra that one band crosses $E=0$ when the parameters obey $|\mu|>2t+t^{\prime 2}/2t$. The line separating the regions where this inequality is satisfied and not satisfied is plotted with red dotted line in Fig. 5(f) and the line coincides with the phase transition from $w=1$ to $w=0$.
V Discussion and Conclusion
The investigation into the behavior of MFs within the framework of the Kitaev model coupled to a one dimensional normal metal (NM) has yielded intriguing results. We began by analyzing the Hamiltonian governing the system, considering parameters such as hopping strengths ($t$), chemical potential ($\mu$), pairing strength ($\Delta$), and coupling between the Kitaev chain (KC) and NM ($t^{\prime}$). Through numerical diagonalization, we explored the bulk spectrum and identified critical points where the gap in the energy spectrum closes. These critical points serve as indicators of phase transitions within the system. Further analysis revealed the presence of zero-energy states, which could potentially harbor MFs. By projecting the Hamiltonian onto different sectors defined by the electron and hole eigenstates, we calculated the winding number- a topological invariant- indicating the presence of MFs. The winding number analysis provided insights into the phase diagram of the system. We find that the exact value of $\Delta$ does not change the phase diagrams as long as it is nonzero. Also, we have taken the chemical potential within the NM to be zero, thereby maintaining the NM in metallic phase. However, by hybridization with the KC, NM can be driven into band insulator phase when the pairing is turned off. An important message of this work is that even when KC and NM are in topologically trivial phase, coupling the two can result in a topological phase.
Ref. DeGottardi et al. (2013a, b) discuss the emergence of Majorna fermions in periodic potentials in a single Kitaev chain. They find that a finite pairing strength that depends on the amplitude of the periodic potential is required to get MFs in the system. The periodic potential on a NM drives it into insulating phase. So, it is interesting to see that MFs can be induced in such systems by a strong enough pairing strength. This is in contrast to our system where the coupled one dimensional system is driven into the topological phase hosting MFs only if the hybrid system is in metallic phase (in absence of pairing) and a nonzero pairing strength on top in one chain is enough to get MFs.
We extended our investigation to include the effect of magnetic flux ($\phi$) threading through the system. The phase diagram exhibited significant changes as $\phi$ varied, with the emergence of new phase transitions and alterations in the configuration of MFs. Notably, for $\phi\geq\pi/2$, the phase diagram simplified to two distinct phases—one with a single pair of MFs and another with none.
These findings underscore the rich behavior of MFs in hybrid systems comprising Kitaev chains and normal metals. The ability to manipulate parameters such as coupling strength and magnetic flux opens avenues for controlling and engineering Majorana fermion states, holding promise for applications in quantum computing and topological quantum information processing.
Acknowledgements.
The author thanks Diptiman Sen for useful comments on the work. The author thanks SERB Core Research grant (CRG/2022/004311) for financial support. The author also thanks funding from University of Hyderabad Institute of Eminence PDF.